# Non-Convex Self-Concordant Functions: Practical Algorithms and Complexity Analysis

### v1

@@ 3.2 Adaptive Regularization Method (ARM)
The key advantage of the ARM over the trust region (TR) method and [Cartis-2022-Evaluation, Algorithm 2.4.1] is to compute a better direction $d_{j}$ and construct a better model $m_{j}(\cdot)$ by leveraging the special self-concordance structure. Indeed, the TR method computes $d_{j}$ based on a local model $m_{j}(d)=f(x_{j})+\nabla f(x_{j})^{\top}d+\frac{1}{2}d^{\top}B_{j}d$ where $B_{j}$ is symmetric such that $f(x_{j}+d_{j})<f(x_{j})$. The corresponding ratio of reduction is $r_{j}=\frac{f(x_{j})-f(x_{j}+d_{j})}{m_{j}(0)-m_{j}(d_{j})}$. The TR method either accepts or rejects but always scales the radius based on $r_{j}$. In [Cartis-2022-Evaluation, Algorithm 2.4.1], a local model $m_{j}(d)=f(x_{j})+\nabla f(x_{j})^{\top}d+\frac{1}{2}\sigma_{j}\|d\|^{2}$ is constructed and $d_{j}=\mathop{\rm argmin}_{d}m_{j}(d)$. All of other steps are similar to that used in the ARM. Another advantage of the ARM is its ability to handle the constrained case where $\textbf{dom}(f)\neq\mathbb{R}^{n}$. It rejects when $x_{j}+d_{j}\notin\textbf{dom}(f)$, while maintaining convergence guarantees.
4 Convergence Analysis
We provide the convergence analysis for the RNM and ARM algorithms proposed in Section 3 by proving all theorems.
4.1 Technical Lemmas
We define a set of successful iterations up to $k$ by $\mathcal{S}_{k}=\{j:0\leq j\leq k,r_{j}\geq\eta_{1}\}$ [^2] We define $r_{i}=0$ if Algorithm 1 terminates when $j\leq i$., unsuccessful iterations up to $k$ by $\mathcal{U}_{k}=\{0,\dots,k\}\backslash\mathcal{S}_{k}$, and all successful iterations by $\mathcal{S}=\bigcup_{k=0}^{\infty}\mathcal{S}_{k}$.
Lemma 4.1.
Suppose that $f$ is bounded below and continuously differentiable. If there exists $\sigma_{\max}>0$ such that $\sigma_{j}\leq\sigma_{\max}$ for all $j\geq 0$ and $\tilde{\epsilon}>0$ such that $f(x_{j})-m_{j}(t_{j})\geq\tilde{\epsilon}$ for all $j\in\mathcal{S}$, Algorithm 1 terminates within $O(\tilde{\epsilon}^{-1})$ iterations.
Proof.
If Algorithm 1 does not terminate up to $k$, we have $f(x_{0})-\inf_{z\in\mathbb{R}^{n}}f(z)\geq\sum_{j\in\mathcal{S}_{k}}f(x_{j})-f(x_{j+1})\geq\sum_{j\in\mathcal{S}_{k}}\eta_{1}\tilde{\epsilon}=\eta_{1}|\mathcal{S}_{k}|\tilde{\epsilon}$. This implies that $|\mathcal{S}_{k}|=O(\tilde{\epsilon}^{-1})$. Then, we see from Algorithm 1 that $\gamma_{1}\sigma_{j}\leq\max(\gamma_{1}\sigma_{j},\sigma_{\min})\leq\sigma_{j+1}$ for $j\in\mathcal{S}_{k}$ and $\gamma_{2}\sigma_{j}\leq\sigma_{j+1}$ for $j\in\mathcal{U}_{k}$. Thus, $\sigma_{0}\gamma_{1}^{|\mathcal{S}_{k}|}\gamma_{2}^{|\mathcal{U}_{k}|}\leq\sigma_{k}$. Since $\sigma_{k}\leq\sigma_{\max}$, we have $|\mathcal{S}_{k}|\log(\gamma_{1})+|\mathcal{U}_{k}|\log(\gamma_{2})\leq\log(\sigma_{\max}\sigma_{0}^{-1})$. This together with $k=|\mathcal{S}_{k}|+|\mathcal{U}_{k}|$ yields $k\leq|\mathcal{S}_{k}|\left(1-\tfrac{\log(\gamma_{1})}{\log(\gamma_{2})}\right)+\tfrac{1}{\log(\gamma_{2})}\log(\frac{\sigma_{\max}}{\sigma_{0}})$. Since $|S_{k}|=O(\tilde{\epsilon}^{-1})$, we have $k=O(\tilde{\epsilon}^{-1})$. ∎
Lemma 4.2.
Suppose that Assumption 3.1 holds and Eq. (9) is satisfied. Let $m_{j}$ be defined in Eq. (10), and the termination condition be $\frac{\nabla f(x_{j})^{\top}d_{j}}{\sqrt{\nabla^{2}(f+\sigma_{j}F)(x_{j})[d_{j},d_{j}]}}\leq\epsilon$. Then, we have $\sigma_{j}\leq\max(\gamma_{3}\bar{\sigma},\sigma_{0},\gamma_{3})$ for all $j$ and $f(x_{j})-m_{j}(t_{j})\geq\frac{\epsilon^{2}}{2(1+\Gamma_{f}(x_{0}))}$ for all $j\in\mathcal{S}$.
Proof.
We prove by contradiction. Suppose that there exists $j$ such that $\sigma_{j+1}>\max(\gamma_{3}\bar{\sigma},\sigma_{0},\gamma_{3})\geq\sigma_{j}$. Then, we have $\sigma_{j}\geq\frac{\sigma_{j+1}}{\gamma_{3}}\geq\max(\bar{\sigma},1)$. We let $\rho_{j}=-\nabla f(x_{j})^{\top}d_{j}$, $\delta_{j}=\nabla^{2}f(x_{j})[d_{j},d_{j}]$, and $\Delta_{j}=\nabla^{2}F(x_{j})[d_{j},d_{j}]$. We obtain from Assumption 3.1 that $\rho_{j}\geq 0$ and $\delta_{j}+\sigma_{j}\Delta_{j}\geq\delta_{j}+\Delta_{j}\geq 0$. We also have
$f(x_{j})-m_{j}(t_{j})=\rho_{j}t_{j}-\kappa^{-2}\omega_{\star}(\kappa t_{j}\sqrt{\delta_{j}+\sigma_{j}\Delta_{j}}).$
By Proposition 2.6, we have $x_{j}+t_{j}d_{j}\in\textbf{dom}(f)$ and
$f(x_{j})-f(x_{j}+t_{j}d_{j})\geq\rho_{j}t_{j}-\kappa^{-2}\omega_{\star}(\kappa t_{j}\sqrt{\delta_{j}+\Delta_{j}}).$
Since $\omega_{\star}(\cdot)$ is increasing, we have $r_{j}\geq 1>\eta_{2}$ and thus $\sigma_{j+1}\leq\sigma_{j}\leq\max(\gamma_{3}\bar{\sigma},\sigma_{0},\gamma_{3})$. This yields a contradiction. Thus, $\sigma_{j}\leq\max(\gamma_{3}\bar{\sigma},\sigma_{0},\gamma_{3})$ for all $j$.
For all $j\in\mathcal{S}$, we provide a lower bound of $f(x_{j})-m_{j}(t_{j})$. Indeed, by combining Eq. (11) with Eq. (10), we have
$f(x_{j})-m_{j}(t_{j})\geq\tfrac{(\nabla f(x_{j})^{\top}d_{j})^{2}}{2(1+\Gamma_{f}(x_{0}))\nabla^{2}(f+\sigma_{j}F)[d_{j},d_{j}]}>\tfrac{\epsilon^{2}}{2(1+\Gamma_{f}(x_{0}))}.$
This completes the proof. ∎
Lemma 4.3.
Suppose that Assumption 3.1 holds with a $\kappa_{F}$-self-concordant $F$. Let $d_{j}$ and $m_{j}$ be defined in Eq. (17) and Eq. (18), and the termination condition be $\nu_{f,\sigma_{j}F}(x_{j})\leq\epsilon_{g}$ and $\lambda_{\min}(\nabla^{2}f(x_{j}))\geq-\sigma_{j}\sqrt{\epsilon_{H}}\nabla^{2}F(x_{j})[v_{j},v_{j}]$. Then, we have $\sigma_{j}\leq\max(\gamma_{3},\sigma_{0})$ for all $j$ and $f(x_{j})-m_{j}(t_{j})\geq\min\left(\frac{\epsilon_{H}^{1.5}}{6(\kappa_{f}+\kappa_{F})^{2}},\frac{\epsilon_{g}^{2}}{2(1+\Gamma_{f}(x_{0}))}\right)$ for all $j\in\mathcal{S}$.
Proof.
We prove by contradiction. Suppose that there exists $j$ such that $\sigma_{j+1}>\max(\sigma_{0},\gamma_{3})\geq\sigma_{j}$. Then, we have $\sigma_{j}\geq\frac{\sigma_{j+1}}{\gamma_{3}}\geq 1$. Using the same idea proving Lemma 4.2, we have $\rho_{j}\geq 0$ and $\delta_{j}+\sigma_{j}\Delta_{j}\geq\delta_{j}+\Delta_{j}\geq 0$. We divide the remaining proof into two cases: (i) $\lambda_{\min}(\nabla^{2}f(x_{j}))\geq-\lambda_{\mathrm{nc}}$; and (ii) $\lambda_{\min}(\nabla^{2}f(x_{j}))<-\lambda_{\mathrm{nc}}$.
For the first case, we have
$f(x_{j})-m_{j}(t_{j})=\rho_{j}t_{j}-\kappa^{-2}\omega_{\star}(\kappa t_{j}\sqrt{\delta_{j}+\sigma_{j}\Delta_{j}}).$
By Proposition 2.6, we have $x_{j}+t_{j}d_{j}\in\textbf{dom}(f)$ and
$f(x_{j})-f(x_{j}+t_{j}d_{j})\geq\rho_{j}t_{j}-\kappa^{-2}\omega_{\star}(\kappa t_{j}\sqrt{\delta_{j}+\Delta_{j}}).$
Since $\omega_{\star}(\cdot)$ is increasing, we have $r_{j}\geq 1>\eta_{2}$ and then $\sigma_{j+1}\leq\sigma_{j}\leq\max(\sigma_{0},\gamma_{3})$. This yields the contradiction.
For the second case, we have $\delta_{j}=\lambda_{\min}(\nabla^{2}f(x_{j}))<0$ and
$f(x_{j})-m_{j}(t_{j})=-\kappa^{-2}\omega_{\star}(\kappa t_{j}\sqrt{\delta_{j}+\sigma_{j}\Delta_{j}})+\kappa_{F}^{-2}\omega(\kappa_{F}t_{j}\sqrt{\sigma_{j}\Delta_{j}}).$
By Proposition 2.6, we have $x_{j}+t_{j}d_{j}\in\textbf{dom}(f)$ and
$f(x_{j})-f(x_{j}+t_{j}d_{j})\geq-\kappa^{-2}\omega_{\star}(\kappa t_{j}\sqrt{\delta_{j}+\Delta_{j}})+\kappa_{F}^{-2}\omega(\kappa_{F}t_{j}\sqrt{\Delta_{j}}).$
Defining $g(\sigma)=-\kappa^{-2}\omega_{\star}(\kappa t_{j}\sqrt{\delta_{j}+\sigma\Delta_{j}})+\kappa_{F}^{-2}\omega(\kappa_{F}t_{j}\sqrt{\sigma\Delta_{j}})$, we have
$g^{\prime}(\sigma)=\tfrac{t_{j}^{2}\Delta_{j}}{2}\left(\tfrac{-1}{1-\kappa t_{j}\sqrt{\delta_{j}+\sigma\Delta_{j}}}+\tfrac{1}{1+\kappa_{F}t_{j}\sqrt{\sigma\Delta_{j}}}\right)\leq 0.$
Thus, $f(x_{j})-m_{j}(t_{j})=g(\sigma_{j})\leq g(1)\leq f(x_{j})-f(x_{j}+t_{j}d_{j})$, which implies $r_{j}\geq 1>\eta_{2}$. This yields a contradiction.
Since Algorithm 1 does not terminate at $x_{j}$ for $j\in\mathcal{S}$, we have $\nu_{f,\sigma_{j}F}(x_{j})>\epsilon_{g}$ or $\lambda_{\min}(\nabla^{2}f(x_{j}))<-\lambda_{\mathrm{nc}}$. Indeed, if $\lambda_{\min}(\nabla^{2}f(x_{j}))\geq-\lambda_{\mathrm{nc}}$ and $\nu_{f,\sigma_{j}F}(x_{j})>\epsilon_{g}$, we obtain from $x-\log(1+x)\geq\frac{x^{2}}{2(1+\Gamma_{f}(x_{0}))}$ for $0\leq x\leq\Gamma_{f}(x_{0})$ that
$f(x_{j})-m_{j}(t_{j})\geq\tfrac{\epsilon_{g}^{2}}{2(1+\Gamma_{f}(x_{0}))}.$
Otherwise, we consider the case of $\lambda_{\min}(\nabla^{2}f(x_{j}))<-\lambda_{\mathrm{nc}}$. For simplicity, we define $\gamma=\frac{\kappa_{F}}{\kappa}$ and $z=-\frac{\delta_{j}}{\sigma_{j}\Delta_{j}}$. Then, Eq. (18) implies
$f(x_{j})-m_{j}(t_{j})=\kappa^{-2}\left(\gamma^{-2}\omega\left(\tfrac{\gamma z}{\gamma(1-z)+\sqrt{1-z}}\right)-\omega_{\star}\left(\tfrac{z}{\gamma\sqrt{1-z}+1}\right)\right)\geq\kappa^{-2}g_{\gamma}(z),$
where
$g_{\gamma}(z)=\tfrac{z}{\gamma\sqrt{1-z}}+\log\left(1-\tfrac{z}{\gamma\sqrt{(1-z)}+1}\right)-\tfrac{1}{\gamma^{2}}\log\left(1+\tfrac{\gamma z}{\gamma(1-z)+\sqrt{1-z}}\right).$
We also have $g_{\gamma}(0)=0$ and $g_{\gamma}^{\prime}(z)\geq\frac{z^{2}}{2(\gamma+1)^{2}}$. This implies $g_{\gamma}(z)\geq\frac{z^{3}}{6(\gamma+1)^{2}}$ and
$f(x_{j})-m_{j}(t_{j})\geq\tfrac{z^{3}}{6(\kappa+\kappa_{F})^{2}}\geq\tfrac{\epsilon_{H}^{1.5}}{6(\kappa_{f}+\kappa_{F})^{2}}.$
This completes the proof. ∎
4.2 Main Results
We are ready to provide the proofs for all theorems.
Proof of Theorem 3.2.
Proposition 2.6 guarantees that $x_{j}\in\textbf{dom}(f)$, $f(x_{j})\leq f(x_{0})$ for all $j\geq 0$ and $f(x_{j})-f(x_{j+1})\geq\frac{(\nu_{f,F}(x_{j}))^{2}}{2(1+\Gamma_{f}(x_{0}))}$. Summing this inequality over $j=0,1,2,\ldots$ yields
$f(x_{0})-\inf_{x\in\mathbb{R}^{n}}f(x)\geq\sum_{i=0}^{k}f(x_{j})-f(x_{j+1})\geq\sum_{i=0}^{k}\tfrac{\nu_{f,F}(x_{j})^{2}}{2(1+\Gamma_{f}(x_{0}))}\geq\tfrac{(k+1)(\min_{0\leq j\leq k}\nu_{f,F}(x_{j}))^{2}}{2(1+\Gamma_{f}(x_{0}))}.$
This completes the proof. ∎
Proof of Theorem 3.3.
We denote by $U$ the neighborhood of $x^{\star}$ where the Polyak-Łojasiewicz condition holds. There exists $L>0$ such that $\|\nabla^{2}(f+F)(x)\|_{\mathrm{op}}\leq L$ for all $x\in U$. By [Absil-2005-Convergence, Proposition 3.3 and Theorem 3.4], we have $x_{j}\to x_{\star}$ if there exist $c_{1},c_{2}>0$ such that $f(x_{j})-f(x_{j+1})\geq c_{1}\|\nabla f(x_{j})\|^{2}$ and $\|\nabla f(x_{j})\|\geq c_{2}\|x_{j}-x_{j+1}\|$ for all $x_{j},x_{j+1}\in U$. Indeed, by Proposition 2.6 and using the fact that $f(x_{j})\leq f(x_{0})$ and $\|\nabla^{2}(f+F)(x_{j})\|_{\mathrm{op}}\leq L$ for all $j$, we have
In addition, we have $\|\nabla f(x_{j})\|\geq\alpha\|x_{j+1}-x_{j}\|$. Putting these pieces together yields the desired result.
It remains to prove the linear convergence of function values. As $x_{j}\to x^{\star}$, there exists $K\in\mathbb{N}$ such that $x_{j}\in U$ for all $j\geq K$. Applying Eq. (20) and the Polyak-Łojasiewicz inequality yields $f(x_{j})-f(x_{j+1})\geq\frac{\mu}{(1+\Gamma_{f}(x_{0}))L}(f(x_{j})-f(x^{\star}))$ for all $j\geq K$. This implies
$f(x_{j+1})-f(x^{\star})\leq\left(1-\tfrac{\mu}{(1+\Gamma_{f}(x_{0}))L}\right)(f(x_{j})-f(x^{\star})).$
which implies that $f(x_{j})-f(x^{\star})$ converges $Q$-linearly to 0. ∎
Proof of Theorem 3.5.
Suppose that Assumption 3.1 holds, Lemma 4.2 implies that Lemma 4.1 holds with $\tilde{\epsilon}=\frac{\epsilon^{2}}{2(1+\Gamma_{f}(x_{0}))}$ and $\sigma_{\max}=\max(\gamma_{3}\bar{\sigma},\sigma_{0},\gamma_{3})$. Thus, Algorithm 1 terminates within $O(\tilde{\epsilon}^{-1})=O(\epsilon^{-2})$ iterations. ∎
Proof of Corollary 3.6.
It follows from Theorem 3.5 that within $O(\epsilon^{-2})$ iterations, we can output $x_{j}$ satisfying
$\tfrac{\sqrt{\nabla f(x_{j})^{\top}H_{j}\nabla f(x_{j})}}{\Lambda}\leq\tfrac{\nabla f(x_{j})^{\top}d_{j}}{\sqrt{\nabla^{2}(f+\sigma_{j}F)[d_{j},d_{j}]}}\leq\tfrac{\epsilon}{\Lambda}.$
This completes the proof. ∎
Proof of Corollary 3.7.
The direction $d_{j}$ and the model $m_{j}$ given by Eq. (14) and Eq. (15), and the termination condition given by Eq. (12) satisfy the conditions of Theorem 3.5 with $\bar{\sigma}=1$. This completes the proof.∎
Proof of Theorem 3.8.
Suppose that Assumption 3.1 holds, Lemma 4.3 implies that the conditions of Lemma 4.1 hold with $\tilde{\epsilon}=\min\left(\frac{\epsilon_{H}^{1.5}}{6(\kappa_{f}+\kappa_{F})^{2}},\frac{\epsilon_{g}^{2}}{2(1+\Gamma_{f}(x_{0}))}\right)$ and $\sigma_{\max}=\max(\sigma_{0},\gamma_{3})$. Thus, Algorithm 1 terminates within $O(\tilde{\epsilon}^{-1})=O(\epsilon_{g}^{-2}+\epsilon_{H}^{-1.5})$ iterations. This completes the proof. ∎
@@ 5 Experiments
We conduct experiments on nonnegative matrix factorization and training convolutional neural networks to evaluate our proposed methods.
@@
We denote $\mathcal{U}[a,b]$ as the uniform distribution on $[a,b]$. For the KL divergence, we generate the data matrix $Z=\hat{X}\hat{Y}+0.01\hat{Z}$, where $\hat{X}\in\mathbb{R}^{m\times r}$, $\hat{Y}\in\mathbb{R}^{r\times n}$, $\hat{Z}\in\mathbb{R}^{m\times n}$ and their entries are sampled i.i.d. from $\mathcal{U}[0,1]$. For the MSE loss, we generate $M=\hat{X}\hat{Y}$, where $\hat{X}\in\mathbb{R}^{m\times r}$, $\hat{Y}\in\mathbb{R}^{r\times n}$ and their entries are sampled i.i.d. from $\mathcal{U}[0,1]$. We compute the singular value decomposition of $M=U\operatorname{diag}(\sigma_{1},\dots,\sigma_{r},0,\dots,0)V^{\top}$ and construct the data matrix $Z=U\mathrm{diag}(\sigma_{1},\dots,\sigma_{r},\sigma_{r+1},\dots,\sigma_{\min(m,n)})V^{\top}$, where $\sigma_{r+1},\dots,\sigma_{\min(m,n)}$ are sampled i.i.d. from $\mathcal{U}[0,0.1\sigma_{r}]$. The optimal function value is given by $\frac{1}{2mn}\|Z-\hat{X}\hat{Y}\|_{F}^{2}$. Thus, the optimality gap is defined as the difference between the current function value and the optimal function value. Throughout, we set $m=100$, $n=20$, and $r=10$. We consider the case where the initialization of $X$ and $Y$ are independent, with each entry drawn i.i.d. from $\mathcal{U}[0,1]$.
We compare the performance of Algorithm 1 (with configurations in Eq. (14) and Eq. (15)) against L-BFGS-B and trust-constr methods from scipy [Virtanen-2020-Scipy]. We also implement several configurations of ARC [Cartis-2011-Adaptive, Cartis-2011-Adaptive2] to illustrate the advantages of our method over cubic regularization method. Since ARC is designed for unconstrained optimization problems, we define $f(X,Y)=+\infty$ whenever $X\notin\mathbb{R}^{m\times r}_{+}$ or $Y\notin\mathbb{R}^{r\times n}_{+}$. In this setting, when some entries of $X$ or $Y$ approach zero, ARC tends to stagnate, with the regularization parameter blowing up since the direction drives those entries toward zero. To mitigate this issue, we incorporate a logarithmic barrier into the objective function when running ARC, defined as
$f_{\mu}(X,Y)=f(X,Y)+\mu\left(\sum_{i,j=1}^{m,r}\log X_{ij}\right)+\mu\left(\sum_{i,j=1}^{r,n}\log Y_{ij}\right).$
We report the original loss function $f(X,Y)$ and tune $\mu$ to achieve the best performance. In both Algorithm 1 and ARC, we use $\sigma_{0}=1$, $\eta_{1}=0.01$, $\eta_{2}=0.9$, $\gamma_{1}=0.5$, and $\gamma_{2}=\gamma_{3}=2$, following [Cartis-2022-Evaluation, Section 2.4]. In Algorithm 1, we set $\kappa=1$.
As shown in Figure 1, Algorithm 1 consistently outperforms other methods. For large $\mu$, although ARC converges faster, it fails to reach high accuracy, while for small $\mu$, ARC may fail to converge. These observations align with the strong performance of ARC in unconstrained optimization and suggest that more refined algorithmic designs may be required for constrained optimization problems.
@@
The use of Newton’s methods in large-scale neural network training is prohibitively expensive, as computing and inverting the Hessian is costly. To overcome this limitation, the Kronecker-Factored Approximate Curvature (KFAC) method was introduced in [Martens-2015-Optimizing] as a scalable alternative to Newton’s method and natural gradient descent.
KFAC is a Kronecker-factorized version of a positive definite Fisher Information Matrix (FIM), which is equivalent to the generalized Gauss-Newton matrix (GGN) in certain cases [Martens-2012-Training, Pascanu-2014-Revisiting, Martens-2020-New, Abreu-2025-Potential], to approximate the Hessian. This avoids doing multiple back propagation steps to compute the exact Hessian and additional storage. KFAC approximates the FIM by exploiting the layer-wise structure of feed-forward neural networks. Indeed, we consider a supervised learning problem with data distribution $(x,y)\sim\mathcal{D}$, where $x$ denotes the input and $y$ the target. A feed-forward neural network defines a parametric function $f_{\theta}:\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{y}}$, with $\theta$ consisting of weight matrices and biases across all layers. In particular, we let the $l$-th layer be $s_{l}=W_{l}a_{l}+b_{l}$, $h_{l}=\phi(s_{l})$, where $a_{l}\in\mathbb{R}^{d_{\textnormal{in}}}$ are the input activation nodes to the layer, $W_{l}\in\mathbb{R}^{d_{\textnormal{out}}\times d_{\textnormal{in}}}$ and $b_{l}\in\mathbb{R}^{d_{\textnormal{out}}}$ are the parameters, $s_{l}\in\mathbb{R}^{d_{\textnormal{out}}}$ are the pre-activations, and $h_{l}\in\mathbb{R}^{d_{\textnormal{out}}}$ are the output activation after applying the elementwise transformation $\phi$.
Algorithm 2 KFAC (i) [Martens-2015-Optimizing, Zhang-2019-Three] and KFAC with weak self-concordance (KFAC-WSC) (ii).
@@
2:  Input: If choosing (ii), input $\kappa$: weak self-concordance parameter, default 1; input $\beta$: layerwise leaning rate momentum, default 0.99.
@@
10:    $G_{l}\leftarrow\nabla_{W_{l}}\mathcal{L}_{B}(W)$, $D_{l}\leftarrow(A_{l}+\gamma I)^{-1}G_{l}(S_{l}+\gamma I)^{-1}$, $M_{l}\leftarrow\mu M_{l}+(1-\mu)D_{l}$.
11:    Compute $\lambda_{l}$ according to one of two options: (i) $\lambda_{l}\leftarrow 0$. (ii) $\lambda_{l}^{2}\leftarrow\beta\lambda_{l}^{2}+(1-\beta)\mathrm{vec}(G_{l})^{\top}\mathrm{vec}(D_{l})$.
12:    $W_{l}\leftarrow W_{l}-\frac{\eta_{k}}{1+\kappa\lambda_{l}}M_{l}-\eta_{k}\omega W_{l}$.

### v2

@@ 3.2 Adaptive Regularization Method (ARM)
The key advantage of ARM over the trust region (TR) method and [13, Algorithm 2.4.1] is that ARM computes a better direction $d_{j}$ and constructs a better model $m_{j}(\cdot)$ by leveraging the special self-concordance structure. Indeed, the TR method computes $d_{j}$ based on a local model $m_{j}(d)=f(x_{j})+\nabla f(x_{j})^{\top}d+\frac{1}{2}d^{\top}B_{j}d$, where $B_{j}$ is a symmetric matrix chosen so that $f(x_{j}+d_{j})<f(x_{j})$. The TR method either accepts or rejects the update $d_{j}$ based on the reduction ratio $r_{j}=\frac{f(x_{j})-f(x_{j}+d_{j})}{m_{j}(0)-m_{j}(d_{j})}$, and always scales the TR radius based on $r_{j}$. In [13, Algorithm 2.4.1], a local model $m_{j}(d)=f(x_{j})+\nabla f(x_{j})^{\top}d+\frac{1}{2}\sigma_{j}\|d\|^{2}$ is constructed and $d_{j}=\mathop{\rm argmin}_{d}m_{j}(d)$. All other steps are similar to those in ARM. Another advantage of ARM is its ability to handle the constrained case where $\textbf{dom}(f)\neq\mathbb{R}^{n}$. It rejects taking a step when $x_{j}+d_{j}\notin\textbf{dom}(f)$, while maintaining convergence guarantees.
@@ 4 Experiments
In this section we report the results of two different types of computational experiments. In the first set of tests we compared the performance of Algorithm 2 (ARM) with option 1 for solving nonnegative matrix factorization problems against that of four other algorithms. In the second set of tests, we modified the well-known approximate natural gradient/Gauss-Newton method, KFAC [47] so that it was transformed into an “approximate Gauss-Newton” version of Algorithm 1 (RNM). We then compared its performance to that of the original KFAC algorithm in training four different convolutional-neural network/data-set combinations, to illustrate the use of our approach to nonconvex self-concordance to methods beyond those analyzed in this paper.
In our experiments, we found that RNM performs well for training convolutional neural networks (CNNs) but is less effective for nonnegative matrix factorization (NMF), which can be explained by the different magnitudes of $F$ in these two cases. In our CNN experiments, $F$ is a quadratic with coefficients on the order of $10^{-2}$, which is consistent with the asymmetric Hessian spectra reported in NN training [83], where negative curvature is mild. In these settings, a small regularizer is sufficient to stabilize the updates and capture local curvature. In contrast, the objective function $f$ and its Hessian entries are much larger in scale for NMF. Near a local minimizer, applying the same globally chosen $F$ introduces excessive regularization, which distorts the update direction and ultimately degrades performance.
@@
Let $\mathcal{U}[a,b]$ denote the uniform distribution on $[a,b]$. For the KL divergence formulation, we generate the data matrix $Z=\hat{X}\hat{Y}+0.01\hat{Z}$, where $\hat{X}\in\mathbb{R}^{m\times r}$, $\hat{Y}\in\mathbb{R}^{r\times n}$, $\hat{Z}\in\mathbb{R}^{m\times n}$ and their entries are sampled i.i.d. from $\mathcal{U}[0,1]$. For the MSE loss, we generate $M=\hat{X}\hat{Y}$, where $\hat{X}\in\mathbb{R}^{m\times r}$, $\hat{Y}\in\mathbb{R}^{r\times n}$ and their entries are sampled i.i.d. from $\mathcal{U}[0,1]$. We compute the singular value decomposition of $M=U\operatorname{diag}(\sigma_{1},\dots,\sigma_{r},0,\dots,0)V^{\top}$ and construct the data matrix $Z=U\mathrm{diag}(\sigma_{1},\dots,\sigma_{r},\sigma_{r+1},\dots,\sigma_{\min(m,n)})V^{\top}$, where $\sigma_{r+1},\dots,\sigma_{\min(m,n)}$ are sampled i.i.d. from $\mathcal{U}[0,0.1\sigma_{r}]$. The optimal function value is given by $\frac{1}{2mn}\|Z-\hat{X}\hat{Y}\|_{F}^{2}$. Thus, the optimality gap is defined as the difference between the current function value and the optimal function value. Throughout, we set $m=100$, $n=20$, and $r=10$. We consider the case where the initialization of $X$ and $Y$ are independent, with each entry drawn i.i.d. from $\mathcal{U}[0,1]$.
We test the performance of Algorithm 2 on NMF with both MSE loss and KL divergence. For the MSE loss (Eq. (2)), we set the base function to $F_{\text{MSE}}(X,Y)=(\|X\|_{F}^{2}+\|Y\|_{F}^{2}+1)^{2}-\sum_{i,k}\log(X_{ik})-\sum_{k,j}\log(Y_{kj})$, while for the KL divergence (Eq. (3)), we set it to $F_{\text{KL}}(X,Y)=-\frac{1}{mn}(\sum_{i,k}\log(X_{ik})+\sum_{k,j}\log(Y_{kj}))$. By Section 2, $f_{\text{MSE}}$ in Eq. (2) is $\ell_{1}F_{\text{MSE}}$-based self-concordant for some $\ell_{1}>0$, and $f_{\text{KL}}$ in Eq. (3) is $\ell_{2}F_{\text{KL}}$-based self-concordant for some $\ell_{2}>0$. We implement Algorithm 2 with Option 1 and the direction $d_{j}=-(\nabla^{2}(f+\sigma_{j}F)(x_{j}))^{+}\nabla f(x_{j})$; see also Lemma 3.6, Eq. (11), and Eq. (12).
We compare the performance of Algorithm 2 against IPOPT (from the python package cyipopt) [77], L-BFGS-B, and trust-constr methods from scipy [76]. We also implement COCARC [14], which applies the framework of adaptive cubic regularization [11, 12] to non-convex optimization with convex constraints, to illustrate the advantages of our method over cubic regularization methods. In both Algorithm 2 and COCARC, we set $\sigma_{0}=1$, $\eta_{1}=0.01$, $\eta_{2}=0.9$, $\gamma_{1}=0.5$, and $\gamma_{2}=\gamma_{3}=2$, following [13, Section 2.4]. In Algorithm 2, we set $\kappa=1$.
As shown in Figure 1, Algorithm 2 (ARM) consistently outperforms the other methods to which it is compared on both the MSE (Frobenius norm) (Eq. (2)) and the KL divergence (Eq. (3)) formulations of NMF. Regarding the theoretical convergence guarantees of the compared methods in these NMF settings and more generally, we note the following. (i) Since the NMF objective has unbounded sublevel sets and unbounded second- and third-order derivatives, COCARC [14] is not guaranteed to converge; and its convergence analysis relies on the Lipschitz constant of the Hessian, which is generally difficult to estimate in practice and can lead to conservative step-size choices. (ii) The convergence of L-BFGS-B on non-convex problems is more subtle: while BFGS enjoys global convergence and superlinear local convergence for convex objectives [60], it is not guaranteed to converge to a stationary point in the nonconvex setting, and counterexamples demonstrating cycling behavior are known [50, 20]; positive results exist, to the best of our knowledge, only under modified update rules such as the cautious BFGS strategy of [41]. (iii) For NMF, ARM requires no knowledge of any Lipschitz constant and adaptively tunes the regularization parameter $\sigma_{j}$, while still enjoying the convergence guarantees established in Theorem 3.4.
@@
The use of Newton’s methods in large-scale NN training is prohibitively expensive, as computing and inverting the Hessian is costly. To overcome this limitation, the Kronecker-Factored Approximate Curvature (KFAC) method was introduced in [47] as a scalable alternative to Newton’s method and natural gradient descent. Specifically, KFAC is an approximate natural gradient method that employs a Kronecker-factorized version of a positive semi-definite Fisher Information Matrix (FIM), which is equivalent to the generalized Gauss-Newton (GGN) matrix in certain cases [48, 61, 49, 1], to approximate the Hessian matrix in a regularized Newton method. It avoids the extra work needed to compute the exact Hessian and further greatly reduces the memory needed by approximating the FIM by a block-diagonal matrix based on the layer-wise structure of NNs. We consider a supervised learning problem with data distribution $(x,y)\sim\mathcal{D}$, where $x$ denotes the input and $y$ the target. A feed-forward NN defines a parametric function $f_{\theta}:\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{y}}$, with $\theta$ consisting of weight matrices and biases across all layers. In particular, in the $l$-th layer, $s_{l}=W_{l}a_{l}+b_{l}$, $h_{l}=\phi(s_{l})$, where $a_{l}\in\mathbb{R}^{d_{\textnormal{in}}}$ is the vector of inputs, $W_{l}\in\mathbb{R}^{d_{\textnormal{out}}\times d_{\textnormal{in}}}$ and $b_{l}\in\mathbb{R}^{d_{\textnormal{out}}}$ are the parameters, and $s_{l}\in\mathbb{R}^{d_{\textnormal{out}}}$ and $h_{l}\in\mathbb{R}^{d_{\textnormal{out}}}$ are, respectively, the vectors of pre-activations and outputs (after element-wise activation $\phi$).
Algorithm 3 KFAC (i) [47, 82] and KFAC with weak self-concordance (KFAC-WSC) (ii).
@@
2:Input: If choosing (ii), input $\kappa$: weak self-concordance parameter, default 1; input $\beta$: layerwise learning rate momentum, default 0.99.
@@
10:    $G_{l}\leftarrow\nabla_{W_{l}}\mathcal{L}_{B}(W)$, $D_{l}\leftarrow-(A_{l}+\gamma I)^{-1}G_{l}(S_{l}+\gamma I)^{-1}$, $M_{l}\leftarrow\mu M_{l}+(1-\mu)D_{l}$.
11:    Compute $\lambda_{l}$ according to one of two options: (i) $\lambda_{l}\leftarrow 0$. (ii) $\lambda_{l}^{2}\leftarrow\beta\lambda_{l}^{2}-(1-\beta)\mathrm{vec}(G_{l})^{\top}\mathrm{vec}(D_{l})$.
12:    $W_{l}\leftarrow W_{l}+\frac{\eta_{k}}{1+\kappa\lambda_{l}}M_{l}-\eta_{k}\omega W_{l}$.